Independent random samples were selected from two quantitative populations, with sample sizes, means, and standard deviations given below.
n1= 55, n2= 65, xbar1= 35.5, xbar2= 31.4, s1= 5.7, s2= 3.3
1.) Construct a 95% confidence interval for the difference in the population means (mu1- mu2). (Round your answers to two decimal places)
2.) Find a point estimate for the fifference in the population means.
3.) Calculate a margin of error. (Round your answer to two decimal places)
Answer #1

65 Ans: n= 55, M2 = 65 x = 35.5 = 31.4, S = 5.7 S₂ = 3.3 .) 954 Gonfidence forderal -(0, -3,2)+7 =(35.5 - 31.4) 1.96 573.3 =